# MITO: Enabling Non-Line-of-Sight Perception using Millimeter-waves through Real-World Datasets and Simulation Tools

### v1

@@ 5 Simulating mmWave Images
In this section, we describe our open-source simulation tool which can be used to generate synthetic mmWave images, and evaluate the simulator’s accuracy.
5.1 Simulation Tool
Our simulation tool takes as input a 3D triangle mesh for
the target object. These mesh files can be found in object datasets (e.g., YCB[17]), taken from online repositories of meshes (e.g., 3D warehouse[1] or TurboSquid[10], with millions of mesh files), or custom-generated on smart-phones (e.g., PolyCam iPhone App[6]).
@@ 5.1 Simulation Tool
Figure 6: Simulation Output. Our simulation takes an a) object mesh as input and produces two different outputs assuming different reflection types: b) specular reflections and c) edge scattering. d/e) We then combine these outputs with different weights to simulate different materials.
@@
Table 1: Simulation Accuracy. 3D F-Score for the combined, specular, and edge simulation.
In addition to the mesh of the object, our simulation also takes as input the radar’s locations to simulate measurements from. At a high level, our simulation tool works by estimating the raw signals expected at each of these locations, given reflections off the target object. Then, synthetic images are produced by feeding these simulated signals into the same image processing pipeline used for real world data (Eq. 1).
Estimating the raw signal at a given location follows three steps:
@@
    where $s_{j}(l,v)$ is the jth sample of the reflection from location $l$ to vertex $v$.
@@
After repeating this for each measurement location, we feed the raw signals through our imaging processing pipeline to produce the final synthetic image.
@@
5.1.1 Specular Reflections.
When an object exhibits purely specular reflections, then signal will reflect off each vertex in only one direction. In some cases, the signal is reflected away from the radar’s receivers, and therefore this signal should not be added to $S_{j}$. Specifically, the signal will only be reflected back towards the radar if the surface normal is pointing directly towards the radar’s location. In any other case, the signal will be reflected away from the receiver.
@@
5.1.2 Edge Reflections.
In other cases, the signal experiences scattering effects from the edges of objects. [^7] While there are more complex theories for how an RF signal reflects off edges[3], they are more computationally expensive to simulate, so we make the assumption that edges scatter uniformly. First, we define a vertex as being on an edge of an object when the angle between two of its faces is greater than an angle $\tau_{e}$, and we find the set of all vertices that lie on an edge, $V_{e}$.
@@
Fig. 6c shows an example when simulating only reflections from edges.
5.1.3 Combing Reflections.
In practice, materials rarely exhibit only one type of reflection, but instead some combination of the three. To account for this, our synthetic images can be added together with different weights to simulate different materials:
@@
where $I_{syn}(\alpha_{1},\alpha_{2})$ is the combined synthetic image for weights $\alpha_{1},\alpha_{2}$, and $I_{s},$and $I_{e}$ are the images from the specular, and edge simulation, respectively. Fig. 6d/e shows two examples of combining simulations with different weights representing two different materials.
@@ 5.2 Simulation Accuracy
To evaluate the accuracy of our simulation tool, we use the standard 3D F-Score metric [45] to compare the real-world and simulation point clouds.
@@ 5.2.1 Aligning the Point Clouds
Before we compute the 3D F-Score, we first need to convert both the real-world and simulation images to a point cloud. To do so, we take the 3D images, and create one point in the center of each voxel which has a power above a certain threshold $\tau_{P}$. Formally:
@@
Next, we align the real-world and simulation point clouds using a standard ICP implementation[4]. This is necessary because our mmWave simulation places the object in the center of the image, while our real world experiments may have objects offset from the center.
@@
where $PR$, $RE$, and $F$ are the precision, recall, and F-Score, respectively, and $\mathds{1}$ is an indicator variable. $N_{R}$ ($N_{S}$) is the number of points in the real (synthetic) point cloud. $d(x,P)$ is the distance between point x and its nearest neighbor in point cloud P: ${\footnotesize d(x,P)=\min\limits_{x^{\prime}\in P}||x-x^{\prime}||}$. ${\small\tau_{F}}$ is the F-Score distance threshold.
Since our goal is not to select a single set of weights for each object, we compute this metric across a range of weights $\{\alpha_{1},\alpha_{2}\}$ and choose the combined simulation which produces the best F-Score (e.g., the simulation which best matches the properties of this real-world object).
$F=\max_{(\alpha_{1},\alpha_{2})\in W}F(I_{syn}(\alpha_{1},\alpha_{2}),I)\vspace{-0.05in}$ (11)
where $F(I_{syn},I)$ is the F-Score for a synthetic image $I_{syn}$ and real-world image $I$. W is the set of all weights.
@@ 5.2.3 Results
Table 1 reports the 25th, 50th, and 75th percentile F-Scores for MITO (1st row), specular only (2nd row) and edge only (3rd row). MITO’s simulator achieves a median F-Score of 94%, while the specular-only or edge-only simulators achieve median F-Scores of 90% and 71%, respectively. This demonstrates the value of MITO’s techniques for combining multiple simulation types.
6 Using MITO for Computer Vision Tasks
6.1 Task 1: Object Segmentation
Here, we show how MITO can be used to build an mmWave object segmentation tool that works in LOS and NLOS.
We build on the state-of-the-art segment-anything model[28]. This model takes as input an image to segment and a prompt, either in the form of points in the image or a bounding box. The model then returns one or multiple segmentation masks, and a score associated with the predicted quality of the mask.
Using an off-the-shelf object segmentation network for mmWave images faces two challenges. First, SAM is trained on 2D RGB images, but mmWave images are single-channel, complex-valued, and 3D, and therefore cannot be used in the SAM model directly. Second, the SAM model requires a prompt before returning segmentation masks.  [^8] While there are fully automatic methods using SAM, such as iterating through a grid of points as a prompt, they aim to provide all the masks within an image, and would not be sufficient in our case since they do not identify which mask contain the object of interest.
Here, we describe how we address these challenges at a high-level and provide additional details in the supplementary. First, we convert our 77 GHz mmWave images to 2D RGB images start by averaging the magnitudes along the depth of the 3D image then use the rainbow colormap for colorization as shown in Fig. 7a. Second, we randomly select a subset of the highest power reflection points as prompts, as shown in Fig. 7b. By feeding the resulting RGB image and prompts to the SAM network, we can obtain mmWave segmentation as shown in the example of Fig. 7c.
Figure 7: mmWave Segmentation. To segment mmWave images, we a) roduce a 2D colorized image, b) select prompt points via a mmWave power-based prompter, and c) use SAM to produce high-quality masks
Figure 8: mmWave Segmentation Performance. The precision and recall for all objects, line-of-sight objects, and fully occluded items.
@@ 6.1 Task 1: Object Segmentation
Scenario Accuracy
  Overall 85%
  Line-of-Sight 89%
Fully Occluded 81%
  Specular-Only
Simulation 76%
Edge-Only
Simulation 68%
@@
Table 2: Classifier Accuracy. Accuracy for overall, LOS, NLOS, using only specular, and only edge simulation.
@@
Qualitative Results. Interestingly, in some cases, the depth camera is unable to measure the depth to every point on the object. This is typically the case for very reflective objects (e.g., mirror-like metallic objects) or clear objects (e.g., glass or plastic). In these cases, MITO is able to reconstruct more of the object than the depth camera. We show a qualitative example of this in Fig. 9. In Fig. 9b, the image from the depth camera has holes due to the object’s material reflectivity. In Fig. 9c, MITO’s mmWave image captures the whole object. Finally, Fig. 9d shows our mmWave segmentation is able to segment the entire object.
Quantitative Results. We evaluate the precision and recall of our segmentation across all objects. Fig. 9 plots the median precision (blue) and recall (red), with error bars denoting the 25th/75th percentiles. The figure shows a high median precision of 92.6%, which indicates that the network is not over-estimating the size of the object. Interestingly, the median recall is lower, at 64%; this is primarily because mmWave reflections are specular, as described earlier. The supplementary provides more details and §8 highlights how future work can improve recall via shape completion.
Fig. 9 additionally shows the precision and recall in the case of line-of-sight objects compared to fully occluded objects. We note that the 25th, 50th, and 75th percentile precisions vary by 2.6%, 2.3%, and 0.03%, respectively, when an object goes from LOS to fully occluded. This shows that MITO is able to accurately image and perceive fully occluded objects with a similar performance as in LOS.
6.2 Task 2: Object Classification (Sim2Real)
@@
Next, we show how MITO can be used to build an object classifier. We design a classifier which takes 77GHz mmWave images as input (after applying the segmentation mask from Sec. 6.1), and outputs the estimated class label. We train the classifier entirely on synthetic data, and evaluate it on real-world images in both LOS and NLOS.
@@
One challenge in training from synthetic images is to determine which types of reflections dominate for a given object. Recall from Sec. 5 that our simulation models two different reflection types. However, it is difficult to accurately measure the material/texture properties of all objects to identify how to correctly combine these simulation images. Instead, we introduce a new data augmentation, whereby we randomly choose different weights for each simulation type and sum the weighted images. This allows the network to train on different material properties for each object. We also incorporate standard image augmentations (rotations, translations, erasing, blur). Additional implementation & evaluation details are in the supplementary.
Results. To evaluate the mmWave object classification, we trained the classifier entirely on synthetic images, and tested it on real-world measurements. Table 2 reports the results. Our classifier performs with 85% accuracy. Furthermore, it achieves an accuracy of 89% and 81% for line-of-sight and fully-occluded objects, respectively. This shows that MITO can be used for end applications such as object classification, and we expect this accuracy will improve significantly in future work.
@@
Furthermore, we perform a microbenchmark to evaluate the impact of our different simulation types on the classifier accuracy. We retrain the classifier using only the specular simulation images (without combining them with the edge images). We repeat this process with only the edge images. Table 2 reports the accuracy of these two partial implementations in rows 4 and 5 for the specular and edge images, respectively. We note that these implementations are only able to achieve 76% and 68% each, while using both simulation types together achieved 85% accuracy. This shows the benefit of our new data augmentation leveraging our different simulation types.
@@
These results shows that MITO has the potential to be used for NLOS applications such as confirming packaged orders in closed boxes. Further, by training on synthetic images and evaluating on real-world data, we demonstrate that the simulation is representative of the real-world.
7 Related Work
Related mmWave Datasets. Recognizing the importance of millimeter-wave perception, past research has made initial steps towards millimeter-wave datasets. However, the vast majority of them are focused either on imaging humans [18, 19, 48, 19, 39, 50] for action recognition, outdoor imaging for autonomous driving [14, 16, 23, 29, 36, 53, 37, 51, 47, 24], or imaging of building interiors (walls, hallways)[15, 38]. In the context of higher resolution imaging of objects, the only dataset that we are aware of is focused on guns and knives for TSA and detecting concealed weapons through body scanners for airport security applications [32, 41, 40]. Our work, however, is focused on generating a dataset of a larger set of generic everyday objects for application such as robotic manipulation and perception.
mmWave Image Simulators. Given the growth in millimeter wave devices for networking and sensing, the vast majority of existing millimeter wave simulations have also focused on understanding channel characteristics (e.g., for 5G networks) [12, 34, 43] or for imaging humans for action recognition and HCI applications [49, 54, 52]. In principle, these simulators could be used also for generating datasets of everyday objects; however, existing simulators are extremely expensive, costing tens of thousands of dollars [12] and/or are limited to simulating metallic items and cannot simulate everyday objects with different properties [40]. In contrast, we developed an open-source simulation tool and demonstrated that it generates synthetic data that matches real-world collected data. We believe that this tool combined with advances in vision algorithms can lead to very high accuracy in non line of sight scenarios.
mmWave Perception. Past work has demonstrated the possibility of various computational tasks using millimeter waves such as autonomous vehicle & pedestrian detection[30, 25], concealed weapon detection[41, 42], car completion for autonomous driving [22, 44], or classification of entirely metallic objects[31]. Our work is inspired by the capabilities that have been demonstrated in these systems and aims to open them to a much broader community and new applications through both a dataset & simulation.
8 Discussions, Limitations, & Future Work
8.1 Non-Line-of-Sight Perception Tasks

### v3

@@ 5 MITO Simulator
In this section, we describe our open-source mmWave image simulator, and evaluate the simulator’s accuracy.
5.1 mmWave Simulation
Our simulator takes as input a 3D triangle mesh and radar locations to simulate mmWave signals from. The mesh files can be sourced from object datasets (e.g., YCB [14]), online repositories (e.g., 3D Warehouse, Turbo Squid), or custom-generated on smartphones (e.g., PolyCam iPhone app).
At a high level, our simulator works by estimating the raw signals expected at each of these locations, given reflections off the target object. Then, synthetic images are produced by feeding these simulated signals into the same image processing pipeline used for real world data (Eq. 1).
@@
Figure 6: Simulation Output. Our simulation uses a a) 3D mesh to produce images assuming different reflection types: b) specular & c) edge. d/e) We combine these with different weights to simulate different materials.
@@
Table 1: Simulation 3D F-Score for combined, specular, & edge simulations
Estimating raw signals at a given location is as follows:
@@
    where $s_{j}(l,v)$ is the jth sample of the reflection from location $l$ to vertex $v$ and back.
@@
After repeating this for each measurement location, we feed the raw signals through our imaging processing pipeline to produce the final synthetic image. Note that we can rotate the mesh files to produce synthetic images at various angles.
@@
5.1.1 Specular Reflections
When an object exhibits purely specular reflections, the signal will reflect off each vertex in only one direction. In some cases, the signal is reflected away from the radar’s receivers, and therefore should not be added to $S_{j}$. Specifically, the signal will only be reflected back towards the radar if the surface normal is pointing towards the radar’s location. Otherwise, the signal will be reflected away from the receiver.
@@ 5.1.1 Specular Reflections
5.1.2 Edge Reflections
In other cases, the signal experiences scattering effects from the edges of objects. First, we define a vertex as being on an edge of an object when the angle between two of its faces is greater than an angle $\tau_{e}$, and we find the set of all vertices that lie on an edge, $V_{e}$.
@@ 5.1.2 Edge Reflections
Fig. 6c shows an example when simulating edge reflections.
5.1.3 Combing Reflections
In practice, materials rarely exhibit only one type of reflection, but instead some combination of the two. To account for this, our synthetic images can be added together with different weights to simulate different materials:
@@ 5.1.3 Combing Reflections
where $I_{syn}(\alpha_{1},\alpha_{2})$ is the combined synthetic image for weights $\alpha_{1},\alpha_{2}$, and $I_{s}$ and $I_{e}$ are the images from the specular and edge simulations, respectively. Fig. 6d-e shows two examples of combining simulations with different weights representing two different materials.
@@ 5.2 Simulation Accuracy
To evaluate the accuracy of our simulator, we use the standard 3D F-score metric [43] to compare the mmWave outputs from real-world and simulation.
@@ 5.2.1 Aligning the Point Clouds
Before we compute the 3D F-score, we first need to convert both the real-world and simulation images to a point cloud. To do so, we apply a standard mmWave point cloud generation procedure[18] by creating one point in the center of each mmWave voxel that has a power above a threshold $\tau_{P}$:
@@
Next, we align the real-world and simulation point clouds using a standard iterative closest point implementation [2]. This is necessary because our mmWave simulation places the object in the center of the image, while our real world experiments may have objects offset from the center.
@@
where $PR$, $RE$, and $F$ are the precision, recall, and F-score. $\mathds{1}$ is an indicator variable. $N_{R}$ ($N_{S}$) is the number of points in the real (synthetic) point cloud. ${\small\tau_{F}}$ is the F-score distance threshold. $d(x,P)$ is the distance from point x to its nearest neighbor in cloud P: ${\footnotesize d(x,P)=\min\limits_{x^{\prime}\in P}||x-x^{\prime}||}$.
Since our goal is not to select a single set of weights for each object, but to allow the simulation to represent many different objects, we compute this metric across a range of weights $\{\alpha_{1},\alpha_{2}\}$ and choose the combined simulation which produces the best F-score (e.g., the simulation which best matches the properties of this real-world object).
$F=\max_{(\alpha_{1},\alpha_{2})\in W}F(I_{syn}(\alpha_{1},\alpha_{2}),I)\vspace{-0.1pt}$ (11)
where $F(I_{syn},I)$ is the F-score for a synthetic image $I_{syn}$ and real-world image $I$, and $W$ is the set of all weights.
@@ 5.2.3 Results
Figure 7: Segmentation a)Colorize 2D image b)Select prompt points c)Segment with SAM
LOS NLOS
Input Pre. Recall F1 IoU Acc. Pre. Recall F1 IoU Acc.
  Camera 99.7% 99.5% 99.4% 98.9% 99.9% 3.6% 100%* 7.0% 3.6% 14.0%
24 GHz 57.9% 50.1% 37.3% 22.9% 93.7% 53.0% 38.7% 29.1% 17.0% 92.1%
77 GHz 98.2% 56.4% 71.4% 55.5% 97.9% 95.4% 57.7% 70.2% 54.1% 97.7%
Multi
Spectral 97.1% 62.6% 74.4% 59.3% 98.0% 95.1% 63.1% 70.8% 54.9% 97.6%
Table 2: Segmentation Accuracy. Precision (Pre), Recall, F1 Score, IoU, and Accuracy (Acc) when segmenting camera, 24 GHz, 77 GHz, and multi-spectral images, in both LOS and NLOS scenarios.
@@
Table 1 reports the 25th, 50th, and 75th percentile F-scores for MITO (1st row), specular only (2nd row) and edge only (3rd row). MITO’s simulator achieves a median F-score of 94%, while the specular-only or edge-only simulators achieve median F-scores of 90% and 71%, respectively. An even larger improvement can be seen in the 25th percentile (85% vs 68% and 49%). This shows the value of MITO’s techniques for combining multiple simulation types.
6 Non-Line-of-Sight Applications
In this section, we evaluate the ability of our dataset and simulator to enable and establish baselines for NLOS perception for diverse, everyday objects. We describe our experimental setup and analyze the evaluation metrics for NLOS segmentation and classification.
Such applications would be useful in a variety of areas, such as confirming customer orders are properly packaged in e-commerce warehouses, detecting broken objects during shipping, and efficiently finding objects for robotic search.
6.1 Object Segmentation
We show how MITO can be used to build a mmWave object segmentation tool that works in LOS and NLOS.
6.1.1 Setup
We build on the state-of-the-art segment-anything model (SAM) [27], which takes as input an image to segment and a prompt, either in the form of points in the image or a bounding box. The model returns one or multiple masks, and a score associated with the predicted mask quality.
Directly using an off-the-shelf object segmentation network for mmWave images faces two challenges. First, SAM is trained on 2D RGB images, but mmWave images are single-channel, complex-valued, and 3D. Second, the SAM model requires a prompt before returning segmentation masks. We note that while there are fully automatic methods using SAM (e.g., iterating through a grid of points as a prompt)[27], they aim to provide all masks within an image, and would not be applicable in our case since they do not identify which mask contains the object of interest.
To convert our mmWave images to 2D RGB images, we start by averaging the magnitudes along the depth of the 3D image, and use a rainbow colormap for colorization, as shown in Fig. 7a. Second, we randomly select a subset of the highest power reflection points as prompts, as shown in Fig. 7b. By feeding the resulting RGB image and prompts to the SAM network, we can obtain mmWave segmentationm as shown in Fig. 7c. We apply the same approach for segmenting both 24 GHz and 77 GHz images. Additional details can be found in the supplementary.
Multi-Spectral Images. Beyond segmenting 77 GHz and 24 GHz images separately, we also propose a method for fusing the two images to improve overall segmentation accuracy. To do so, we combine the images during prompt point selection: instead of using a power threshold based only on a single image, the points must exceed thresholds in both 24 GHz and 77 GHz images. Then, we use these points and the 77 GHz colorized image as input to the SAM network. This allows the 24 GHz image to act as a filter, removing the impact of noise (e.g., from cardboard occluders) on the point selection and improving segmentation quality.
6.1.2 Experiments
Baseline. We compare our segmentation performance to an RGB-D camera baseline. We use the camera image from the experiment directly as input to the SAM model. To select prompt points, we use the depth image to automatically filter out pixels greater than a certain depth (i.e., the background), and randomly select prompt points from the remaining pixels. In the LOS scenarios, this corresponds to points on the object. In NLOS scenarios, this corresponds to points on the occluding object (i.e., the cardboard).
@@ 6.1.2 Experiments
Training Data Accuracy
  Edge Specular LOS NLOS Overall
  X 61.1% 75.0% 67.6%
X 77.7% 75.0% 76.4%
X X 88.8% 81.2% 85.2%
@@
Table 3: Classifier Accuracy for LOS, NLOS,& overall using specular, edge,& combined simulations for training.
Quantitative Results. Tab. 2 shows the median Precision, Recall, F-score, Intersection-over-Union (IoU), and Accuracy when using the camera, 24 GHz, 77 GHz, and multi-spectral (24 GHz & 77 GHz) images for both LOS and NLOS scenarios. In LOS, segmenting camera images achieves a median precision, recall and F-score of 99.7%, 99.5%, and 99.4%, respectively. However, in NLOS, the camera only achieves 3.6% precision and 7.0% F-score, which is expected since cameras (and other visible-light-based sensors) are unable to operate through occlusions. We note that the camera still achieves 100% recall since SAM segments the cardboard occluder, covering the full object.
In contrast, the various mmWave image inputs achieve better performance NLOS scenarios, demonstrating how MITO enables NLOS object segmentation. In particular, the F-score of NLOS for 24 GHz, 77 GHz, and multi-spectral exceeds that of the camera. Notably, multi-spectral images achieve the highest performance. When comparing LOS and NLOS, multi-spectral images only experience a 2.0%, 0.5%, and 3.6% change in precision, recall, and F-score, respectively. This demonstrates that MITO is able to accurately image and perceive fully occluded objects with a similar performance as in LOS.
Interestingly, the recall when segmenting mmWave images is lower (63.1%) than the precision (95.1%). This is primarily because mmWaves are specular, as described earlier, preventing certain portions of the object from appearing in the image. We provide a more detailed discussion on this phenomenon in the supplementary and Sec. 6.3 highlight how future work can improve recall via shape completion.
The IoU and accuracy reported in Table 2 show similar trends. Notably, 24 GHz, 77 GHz, and multi-spectral all outperform the camera in NLOS. We also note that the accuracy is high because it is dominated by the background.
@@
Qualitative Results. In some cases, the depth camera is unable to measure the depth to every point on the object. This is typically the case for very reflective objects (e.g., mirror-like metallic objects) or clear objects (e.g., glass or plastic). In these cases, MITO is able to reconstruct more of the object than the depth camera, as shown in Fig. 8. In Fig. 8b, the image from the depth camera has holes due to the object’s material reflectivity, whereas in Fig. 8c, MITO’s mmWave image captures the whole object. Fig. 8d shows our 77 GHz mmWave segmentation masks the entire object.
6.2 Sim2Real Shape Classification
Next, we demonstrate how MITO can be used to classify objects in NLOS based on their geometry.
6.2.1 Setup
@@
We design a classifier which takes 77 GHz mmWave images (after applying the segmentation mask from Sec. 6.1) as input, and outputs the estimated class label. The classifier is a neural network consisting of 5 convolutional layers, followed by 3 fully connected layers, and a softmax activation. We train the classifier entirely on synthetic data, and evaluate on real-world images in both LOS and NLOS. To generate synthetic images for training, we leverage object meshes provided by YCB dataset.
@@
One challenge in training from synthetic images is to determine which types of reflections dominate for a given object. Recall from Sec. 5 that our simulator models two different reflection types. However, it is difficult to accurately measure the material/texture properties of all objects to identify how to correctly combine these simulation images. Instead, we introduce a new data augmentation, where we choose random weights for each simulation type and sum the weighted images. This allows the network to train on different material properties for each object. We also apply standard image augmentations (rotation, translation, masking, blur). More training details are in supplementary.
6.2.2 Experiments
We report the shape classification results on real-world data in LOS and NLOS in Tab. 3. Our classifier performs with an overall accuracy of 85.2%. Specifically, it achieves an accuracy of 88.8% and 81.2% for LOS and NLOS scenarios, respectively, showing that MITO can be used to enable novel capabilities such as NLOS shape classification.
@@ 6.2.2 Experiments
In Tab. 3, we also include results using a classifier trained on either only edge images (i.e., without combining them with the specular images) or only specular images in order to evaluate the impact of our different simulation types on the classifier accuracy. We observe that these models are only able to achieve 67.6% and 76.4%, whereas using both simulation types together achieves 85.2% accuracy. This shows the benefit of our simulator’s different outputs, and our data augmentation technique to combine them.
@@
These results shows that MITO has the potential to be used for novel NLOS applications, such as confirming packaged orders in closed boxes. Furthermore, by training on synthetic images and evaluating on real-world data, we show that our simulator is representative of the real-world.
6.3 Additional NLOS Applications & Extensions